Somerset Corporation makes a special-purpose machine, S4H, used in the textile industry. Somerset has designed the S4H machine for 2018 to be distinct from its competitors. It has been generally regarded as a superior machine. Somerset presents the following data for 2017 and 2018.

2017

2018

Units of S4H produced and sold

5,000

5,500

Selling price

$400

$440

Direct materials (pounds)

15,000

15,375

Direct materials costs per pound

$40

$44

Manufacturing capacity for S4H (units)

10,000

10,000

Conversion costs

$1,000,000

$1,100,000

Conversion costs per unit of capacity

$100

$110

Selling and customer-service capacity (customers)

60

58

Total selling and customer-service costs

$360,000

$362,500

Selling and customer-service capacity cost per customer $6,000                    $6,250

Somerset produces no defective machines, but it wants to reduce direct materials usage per S4H machine in 2017. Conversion costs in each year depend on production capacity defined in terms of S4H units that can be produced, not the actual units produced. Selling and customer-service costs depend on the number of customers that Somerset can support, not the actual number of customers it serves. Somerset has 75 customers in 2017 and 80 customers in 2018.

Required:

  1. Is Somerset’s strategy one of product differentiation or cost leadership? Explain briefly.
  2. Describe briefly key measures that you would include in Somerset’s balanced scorecard and the reasons for doing so.
  3. Calculate the the revenue and cost effect of the price-recovery component, and What is the net effect on operating income as a result of the price-recovery component?
Answer #1

Growth component

Revenue

= (CY units - LY units)*LY Price

= (240 - 225)*45,000

= 675,000 F

Material cost

= (Units required for CY sales - LY units)*LY Price

= (300,000/225*240 - 300,000)*8.25

= 165,000 U

Selling and customer-service costs

= (CY capacity - LY capacity) * LY rate

= (115 - 120) * 10,000

= 50,000 F

Price recovery component

Revenue

(CY price - LY price)*CY units

=(47,000 - 45,000)*240

= 480,000 F

Material cost

=(CY price - LY price)*Units required for CY sales

=(8.50*8.25)*320,000

=80,000 U

Conversion cost

=CY cost - LY cost

= 2,100,000 - 2,062,500

=37,500 U

Selling and customer-service costs

=(CY rate - LY rate)*CY capacity

= (9,800 - 10,000)*115

= 23,000 F

Productivity component

Material cost

=(Units required for CY sales - Units used in CY)*CY price

= (310,000 - 300,000/225*240)*8.50

= 85,000 F

Growth component indicates the effect of growth in volume of the sales on the income.

Price recovery component indicates the effect of the changes in the sales price and cost prices on the income.